The speed of sound in air is 341ms-1 at a temperature of 290K and a pressure of1.00×105Pa. If

the temperature falls to 285K and the pressure of1.02×105Pa, does the speed of sound increase

  or decrease? Justify your statement.

What would be the effect of such a change of atmospheric conditions on the frequency of sound

emitted by an organ pipe?



1
Expert's answer
2021-03-04T11:57:43-0500

The speed of sound would not depend on the pressure in such conditions. At 285K, it will be


"c=331.3\sqrt{1+\frac{\theta}{273.15}}=331.3\sqrt{1+\frac{285-273.15}{273.15}}=\\=338\text{ m/s}."

It will decrease.

Such a change of atmospheric conditions would not cause any changes in the frequency because it does not depend on temperature and pressure.